OpenStudy (anonymous):

solve inequality . 3 is less than or equal to 2x-1<5

OpenStudy (swissgirl):

\(3 \leq 2x-1<5\) Our goal is to get the x isolated meaning separated by itself So our first step would be to add 1 to every part of the inequality \(3+1\leq 2x-1+1<5+5 \) \(4 \leq 2x <6\) Next we will divide every part by 2 \( \large\frac {4} {2} \leq \frac{2} {2}x < \frac {6} {2} \) \(2 \leq x <3\) And this the answer :)